# Evaluating the Utility of Conformal Prediction Sets for AI-Advised Image Labeling

### v1

@@ Appendix B Additional Results
B.1. Shortest Path Length
@@
B.1.1. In-distribution
At a high level, the SP length is, on average, smaller for easy images across conditions. When task images are hard, the SP length is greater for images with large set sizes across conditions.
When task images are easy, the SP length does not appear to be affected by set size—the distributions of expected SP length closely overlay across conditions (cf. Figure 7, (A) and (B)). After marginalizing by set sizes, we find that when participants have access to predictions, even when the provided label is wrong, the incorrect label appears to be closer to the correct label in distance (HPI: ) than those of the baseline condition (HPI: ).
@@ B.1.1. In-distribution
For hard task images with a larger average set size of , all intervals in Figure 7 (D) get wider than those in (C), indicating larger variations. Although SP length increased across conditions, we did not detect noticeable differences among conditions with predictions (HPI: ). Similarly, the baseline participants produce more inaccurate responses with a median SP length of (HPI: ).
In summary, we find that the SP length mirrors accuracy in that prediction display that yields higher accuracy tends to produce a lower SP length. Between accuracy and SP length, a noticeable difference can be detected from task images that are hard with a smaller set size: although participants in the Top-1 and the baseline conditions have similar labeling accuracy, the SP length of Top-1 participants is noticeably lower than that of the baseline. It implies that predictions can help elicit responses closer to the correct label, even if they are wrong. Otherwise, our SP results are mostly consistent with our accuracy results.
B.1.2. Out-of-distribution
For OOD task images, the SP length between easy and hard images becomes more contrasting. For easy images, because predictions are still reliable and accurate, we find that the Top-1 condition tends to produce the lowest SP length of (HPI: ) than Top-10 (; HPI: ) and RAPS (; HPI: ), after marginalizing over small and large set sizes (cf. Figure 7, (E) and (F)).
For hard OOD task images, we see an apparent mirroring effect relative to accuracy. When image stimuli have smaller average set size of shown in Figure 7 (G), Top-1 participants have the lowest accuracy rate and also provide labels that have the highest SP length of (HPI: ). RAPS participants have the highest accuracy for this type of image, produce the lowest SP length of (HPI: ). The Top-10 condition overlaps with RAPS with an SP length of (HPI: ), and both Top-10 and RAPS produce lower SP lengths than Top-1 and the baseline condition (; HPI: ).
@@ B.1.2. Out-of-distribution
However, despite that participants of the prediction set have similar accuracy for hard OOD images with small and large sets, we find SP length is smaller for images with larger set sizes. This implies that a larger set size is helpful for hard OOD images, guiding participants to the labels that are closer to the ground truth.
B.2. Distribution of Willingness-to-Pay
Condition Min. Q1 Q2 Q3 Max. Mean SD RAPS 0 1 1.0 2 8 1.78 1.74 Top-1 0 1 1.0 2 8 1.70 1.56 Top-10 0 1 1.5 2 8 1.84 1.67
Distribution of willingness-to-pay elicited from participants.
B.3. Qualitative Analysis of Strategies
Summarizing the qualitative results per identified open code.

### v6

@@ Appendix B Additional Results
B.1. Qualitative Analysis of Strategies
Figure 10 presents a trellis plot summarizing the qualitative results per identified open code. Each row presents a code category with columns showing a barchart summarizing counts of each unique code from participants’ strategies by treatment conditions.
This figure presents a trellis plot summarizing the qualitative results per identified open code. Each row presents a code category with columns showing a barchart summarizing counts of each unique code from participants’ strategies by treatment conditions.
B.2. Shortest Path Length
@@
B.2.1. In-distribution
At a high level, the SP length is, on average, smaller for easy images across conditions. When task images are hard, the SP length is greater for images with larger set sizes across conditions.
When task images are easy, the SP length does not appear to be affected by set size—the distributions of expected SP length closely overlay across conditions (cf. Figure 7, A and B). After marginalizing by set sizes, we find that when participants have access to predictions, even when the provided label is wrong, the incorrect label appears to be closer to the correct label in the distance (HPI: ) than those of the baseline condition (HPI: ).
@@ B.2.1. In-distribution
For hard task images with a larger average set size of , all intervals in Figure 7 (D) get wider than those in (C), indicating greater variations. Although SP length increased across conditions, we do not detect noticeable differences among conditions with predictions (HPI: ). Similarly, the baseline participants produce more inaccurate responses with a median SP length of (HPI: ).
In summary, we find that the SP length mirrors the accuracy in that prediction display that yields higher accuracy, which tends to produce a lower SP length. Between accuracy and SP length, a noticeable difference can be detected from task images that are hard with a smaller set size: although participants in the Top-1 and the baseline conditions have similar labeling accuracy, the SP length of Top-1 participants is noticeably lower than that of the baseline. This implies that predictions can help elicit responses closer to the correct label, even if they are wrong. Otherwise, our SP results are mostly consistent with our accuracy results.
B.2.2. Out-of-distribution
For OOD task images, the SP length between easy and hard images becomes more contrasting. For easy images, because predictions are still reliable and accurate, we find that the Top-1 condition tends to produce the lowest SP length of (HPI: ) than Top-10 (; HPI: ) and RAPS (; HPI: ), after marginalizing over smaller and larger set sizes (cf. Figure 7, E and F).
For hard OOD task images, we see an apparent mirroring effect relative to accuracy. When image stimuli have a smaller average set size of shown in Figure 7 (G), Top-1 participants have the lowest accuracy rate and also provide labels that have the highest SP length of (HPI: ). RAPS participants have the highest accuracy for this type of image and produce the lowest SP length of (HPI: ). The Top-10 condition overlaps with RAPS with an SP length of (HPI: ), and both Top-10 and RAPS produce lower SP lengths than Top-1 and the baseline condition (; HPI: ).
@@ B.2.2. Out-of-distribution
However, despite that participants of the prediction set have similar accuracy for hard OOD images with smaller and larger sets, we find SP length is smaller for images with larger set sizes. This implies that a larger set size is helpful for hard OOD images, guiding participants to labels closer to the ground truth.
B.3. Distribution of Willingness-to-Pay
Table 10 presents the summary statistics of the elicited willingness-to-pay, while Figure 11 visualizes the distributions using a dot plot.
| Condition | Min. | Q1 | Q2 | Q3 | Max. | Mean | SD |
| RAPS | 0 | 1 | 1.0 | 2 | 8 | 1.78 | 1.74 |
| Top-1 | 0 | 1 | 1.0 | 2 | 8 | 1.70 | 1.56 |
| Top-10 | 0 | 1 | 1.5 | 2 | 8 | 1.84 | 1.67 |
This figure presents the distribution of willingness-to-pay elicited from participants by treatment conditions.